# Resilient Distributed Optimization for Multi-Agent Cyberphysical Systems

### v2

@@ Abstract
This work focuses on the problem of distributed optimization in multi-agent cyberphysical systems, where a legitimate agents’ iterates are influenced both by the values it receives from potentially malicious neighboring agents, and by its own self-serving target function. We develop a new algorithmic and analytical framework to achieve resilience for the class of problems where stochastic values of trust between agents exist and can be exploited. In this case we show that convergence to the true global optimal point can be recovered, both in mean and almost surely, even in the presence of malicious agents. Furthermore, we provide expected convergence rate guarantees in the form of upper bounds on the expected squared distance to the optimal value. Finally, numerical results are presented that validate our analytical convergence guarantees even when the malicious agents compose the majority of agents in the network and where existing methods fail to converge to the optimal nominal points.
@@ 1 Introduction
We are interested in investigating the class of problems where the physicality of the system plays an important role in achieving new possibilities of resilience for these systems. Indeed, the physicality of cyberphysical systems has been shown to provide many new channels of verification and establishing inter-agent trust through watermarking [21], wireless signal characteristics [20, 22], side information [23], and camera or lidar data cross-validation [24]. By exploiting these physical-based measurements, agents can extract additional information about the trustworthiness of their neighbors.
@@ 1.1 Related work
To harm the system, a malicious agent can send falsified data to their legitimate neighbors. If the legitimate agents are unaware that this data comes from malicious neighbors, then the malicious agents will succeed in controlling the system [34, 35, 36, 10, 11, 37, 38, 39]. To combat the harmful effect of an attack, the approach taken in [34, 35, 36] requires the pre-existence of a set of trusted agents such that all other agents (legitimate or malicious) are connected to at least one trusted agent. Nonetheless, this approach is unrealistic when communication is sporadic such as in robotic and ad-hoc networks. The approaches in [10, 11, 40, 37, 38, 39] rely on the agent data values to detect and discard malicious inputs and either assume that all the legitimate agents have at least one common local minimizer [39], have an upper bound on the number of tolerable malicious agents (with a star network having the largest number - a half of the number of agents in the network) [37]. When the number of malicious agents exceeds the tolerable number, the attack succeeds and malicious agents evade detection. In contrast with the existing works, our proposed method provides a significantly stronger resilience to malicious activity by exploiting the physical aspect of the problem, i.e., the wireless medium. Thus, each legitimate agent can learn trustworthy neighbors while optimizing the system objective. Our prior work [41] studies the implications of the agents’ learning ability, with regards to the trustworthiness of their neighbors, on distributed consensus systems. This work considers the more general case of distributed optimization systems where the agent’s goal is to minimize the sum of their local objective function under limited information exchange.
@@ 1.2 Paper contribution
This work extends our earlier paper [1], which formalized the problem of resilient distributed optimization with trust observations and presented an upper bound on the convergence rate in the mean squared sense. This work excluded the proofs for these theoretical results due to space limitations. The key contributions of this work beyond those of [1] are as follows: 1) We present the analysis of all the proofs that were omitted from [1]. 2) We provide additional insights and discussions for the theoretical results that are included in [1]. Specifically, this work includes extensive discussions regarding the assumptions that are used in this paper regarding the statistics of the trust observations and their relaxations. 3) We extend the convergence in the second moment that follows from [1] and establish the almost sure convergence to the true optimal point. Additionally, we derive convergence in the $r$th mean for any $r>0$, both as a result of the almost sure convergence and by definition. These convergence results lead to an additional method to derive upper bounds on the convergence rate of the second moment which we present in this work. 4) We present additional numerical simulations that significantly extend the one-dimensional consensus with constraint setup to higher dimensional constraint optimization setups. Furthermore, we present in the supplementary material a comparison between the numerical simulations and the upper bounds we establish in this work.
@@ 2 Problem Formulation
We study the case where an unknown subset of the agents is malicious and the trustworthy agents are learning which neighbors they can trust. Thus, $\mathbb{V}=\mathcal{L}\cup\mathcal{M}$ where $\mathcal{L}$ is the nonempty and finite set of legitimate agents that execute computational tasks and share their data truthfully, while $\mathcal{M}$ denotes a finite set of agents that are not truthful. The sets $\mathcal{L}$ and $\mathcal{M}$ are just modeling artifacts, and none of the legitimate agents knows if it has malicious neighbors, or not, at any time. Throughout the paper, we will use the subscripts $\mathcal{L}$ and $\mathcal{M}$ to denote the various quantities related to legitimate and malicious agents, respectively.
@@ 2.1 Trust values
This model of inter-agent trust observations has been used in prior works [22, 41]. In general, the trust observations can be extracted from the physicality of the systems in various ways as elaborated in the recent survey [50]. For example, the pioneering works [51, 52] extract trust observation based on the characteristics of the received wireless signals that are used to transfer information between agents. The trust values then capture the likelihood of discrepancies between the agent’s reported location and its true location that is manifested in the signal characteristics, such as power and direction of arrival.
@@
(ii) [Independence of trust observations] The observations $\alpha_{ij}(t)$ are statistically independent for all $t$ with $i\in\mathcal{L}$, $j\in\mathcal{N}_{i}$.
@@
Discussion: For the sake of simplicity of presentation, we assume that the expectation terms $E_{\mathcal{L}}$ and $E_{\mathcal{M}}$ are homogeneous (see Assumption 1(i)). We note that the closed-form analytical results derived in this work can be readily extended to the case where $E_{\mathcal{L}}\triangleq\min_{i\mathcal{L},j\in\mathcal{N}_{i}\cap\mathcal{L}}\liminf_{t\rightarrow\infty}\{\mathbf{E}[\alpha_{ij}(t)]-0.5\}$ and $E_{\mathcal{M}}\triangleq\min_{i\in\mathcal{L},j\in\mathcal{N}_{i}\cap\mathcal{M}}\mathop{\rm limsup}_{t\rightarrow\infty}\{\mathbf{E}[\alpha_{ij}(t)]-0.5\}$ are such that $E_{\mathcal{L}}>0$ and $E_{\mathcal{M}}<0$. Additionally, our analysis can be readily adapted to incorporate heterogeneity across clients and replace the uniform $E_{\mathcal{L}}$ and $E_{\mathcal{M}}$ terms with per link terms, i.e., $E_{ij}=\mathbf{E}[\alpha_{ij}(t)]-0.5$. Additionally, while in this work we derive a first-moment analysis by utilizing Hoeffding’s inequality, one can derive an analogous second-moment analysis by utilizing Bernstein’s inequality. Finally, we include in Section 4 a discussion regarding the relaxation of Assumption 1(ii) to include random processes with memory such as Markov chains and martingales.
@@ 2.2 The update rule of agents
The update rule of legitimate agents
@@
To increase the convergence rate, we allow $\gamma(t)$ and $w_{ij}(t)$ to depend on a parameter $T_{0}\geq 0$ which dictates how many trust observations a legitimate agent collects before it decided to trust one of its neighbors. We specify $\gamma(t)$ and $w_{ij}(t),\>i\in\mathcal{L}$, $j\in\mathcal{N}_{i}\cup\{i\}$, precisely later on in Section 3.
@@
Malicious agents $i\in\mathcal{M}$ choose values in the set $\mathcal{X}$ and they may or may not collaborate in choosing their values. We assume that their actions are not known, and thus we do not model them, but assume the worst-case scenario instead in our analysis. For simplicity of exposition, the dynamic (2.2) captures malicious inputs, where an adversarial agent $i\in\mathcal{M}$ sends all its legitimate neighbors identical copies of its chosen input $x_{i}(t)$ at time $t$. Let us denote by $x_{ij}(t)$ the input of a malicious agent $i$ to a legitimate agent $j$ at time $t$, then $x_{ij_{1}}(t)=x_{ij_{2}}(t)$ for every $j_{1},j_{2}\in\mathcal{L}$. Nonetheless, our analytical results also hold for byzantine inputs where an adversarial agent $i\in\mathcal{M}$ can send its legitimate neighbors different inputs at time $t$. In this case $x_{ij_{1}}(t)$ need not be equal to $x_{ij_{2}}(t)$ for every $j_{1},j_{2}\in\mathcal{L}$.
@@ 2.3 Assumptions on connectivity and objective functions
The subgraph $\mathbb{G}_{\mathcal{L}}$ induced by the legitimate agents is connected.
@@ 3 Algorithm
Next, we present an algorithm that incorporates the legitimate agents’ learning of inter-agent trust values into the dynamic (2.2) through the choice of the time-dependent weights $w_{ij}(t)$. These weights depend on a parameter $T_{0}$ that captures the number of trust measurements a legitimate agent collects before deciding whether or not to trust one of its neighbors. We utilize the parameter $T_{0}$ to enable faster convergence rates of the algorithm, nonetheless $T_{0}$ is mostly used for analysis and it is not strictly necessary. In other words, as we show in Section 5, the algorithm converges to the optimal point $x_{\mathcal{L}}^{\star}$ for any choice of nonnegative integer $T_{0}$, including the special case where $T_{0}=0$. In this case, legitimate agents have no prior trust observations to rely on when they first decide whether to trust their neighbors.
@@ 3.1 The weight matrix sequence
$\beta_{ij}(t)=\sum_{k=0}^{t-1}\left(\alpha_{ij}(k)-0.5\right)\hbox{ \ for $t\geq 1,i\in\mathcal{L},\ j\in\mathcal{N}_{i}$},$ (4)
@@
We define the weight matrix $W(t)$ by choosing its entries $w_{ij}(t)$ as follows: for every $i\in\mathcal{L}$, $j\in\mathcal{N}_{i}$,
$w_{ij}(t)=\begin{cases}\frac{\mathds{1}_{\{t\geq T_{0}\}}}{2\cdot\max\{d_{i}(t),d_{j}(t)\}}&\text{if }j\in\mathcal{N}_{i}(t),\\ 0&\text{if }j\notin\mathcal{N}_{i}(t)\cup\{i\},\\ 1-\displaystyle{\sum_{m\in\mathcal{N}_{i}}w_{im}(t)}&\text{if }j=i.\end{cases}$ (6)
@@
Furthermore, the dependence of the weights $w_{ij}(t)$ on the trust observation history $\beta_{ij}(t)$ comes in through the choice of time-dependent and random trusted neighborhood $\mathcal{N}_{i}(t)$ (see (5)). Consequently, some entries of the matrix $W(t)$ are also random, as seen from (6). The gradients $\nabla f_{i}(c_{i}(t))$ are stochastic due to the randomness of $c_{i}(t)$, however, they are not unbiased as typically assumed in stochastic approximation methods, including [53]. Thus, we cannot readily rely on prior analysis for stochastic approximation methods. However, as we show in our subsequent analysis, the variance of $\|\nabla f_{i}(c_{i}(t))\|$ decays sufficiently fast and allows convergence to the optimal point even in the presence of malicious agents.
@@ 4 Learning the sets of trusted neighbors
We can now conclude that there is a random but finite time $T_{f}$ such that there exists a legitimate agent $i$ which misclassifies the trustworthiness of at least one of its neighbors at time $T_{f}-1$, and all the legitimate agents classify the trustworthiness of their neighbors correctly at each time $t\geq T_{f}$. We refer to the time $T_{f}$ as the “correct classification time”.
@@
$\beta_{ij}(t)$ $<0\text{ for all $t\geq T_{f}$ and all }i\in\mathcal{L},j\in\mathcal{N}_{i}\cap\mathcal{M},$ (8)
and there exists $i\in\mathcal{L}$ such that
@@
$\beta_{ij}(T_{f}-1)$ $\geq 0\text{ for some }j\in\mathcal{N}_{i}\cap\mathcal{M},$ (9)
if $T_{f}\geq 1$.
@@
Furthermore, we define the following upper bound on the probability that a legitimate agent misclassifies one of its legitimate or malicious neighbors, in one of the times after observing $k$ trust values for each of its neighbors:
@@ 5 Asymptotic Convergence to Optimal Point
$\leq\gamma(t-T_{0})\|\nabla f_{i}(c_{i}(t))\|\leq\gamma(t-T_{0})G.$
Let us denote $d_{i,\mathcal{L}}\triangleq|\mathcal{N}_{i}\cap\mathcal{L}|+1$. Next, we define the doubly stochastic matrix $\overline{W}_{\mathcal{L}}\in[0,1]^{|\mathcal{L}|\times|\mathcal{L}|}$ with the entries $[\overline{W}_{\mathcal{L}}]_{i,j}$, for every $i,j\in\mathcal{L}$:
@@
Note that $\overline{W}_{\mathcal{L}}$ is the nominal weight matrix, i.e., the value the weight matrix would take in the absence of malicious agents. Let $\sigma_{2}(A)$ be the second largest singular value of $A$, and denote $\rho_{\mathcal{L}}=\max_{k\geq 1}\sigma_{2}(\overline{W}_{\mathcal{L}}^{k})$. Since $\mathbb{G}_{\mathcal{L}}$ is connected and $\overline{W}_{\mathcal{L}}$ is doubly stochastic, the value $\rho_{\mathcal{L}}<1$ is equal to the second largest eigenvalue modulus of $\overline{W}_{\mathcal{L}}$. Additionally, our analysis holds for the lazy Metropolis weight as well for which by [56, Lemma 2.2], the value $\rho_{\mathcal{L}}$ can be upper bounded by $(1-1/(71|\mathcal{L}|^{2}))$, while [57] improves the constant $1/71$ to $1/8$.
Next, define for $t\geq 0$ the following deterministic dynamic that excludes malicious agents,
@@
Denote for $T\geq 1$, the auxiliary error function which we utilize to characterize the convergence rate:
$\overline{h}(T)\triangleq\frac{G^{2}T}{\mu}+\frac{2G^{2}T}{\mu(1-\rho_{\mathcal{L}})}+\frac{8(\mu+L)G^{2}}{\mu^{2}(1-\rho_{\mathcal{L}})^{2}}\ln\left(\frac{T+2}{2}\right)+\frac{2\eta G}{1-\rho_{\mathcal{L}}}$
$+\frac{2(\mu+L)(\mu\eta+2G)^{2}}{\mu^{2}(1-\rho_{\mathcal{L}})^{2}}+\frac{2G^{2}+4G\eta(\mu+L)}{\mu(1-\rho_{\mathcal{L}})^{3}}+\frac{G^{2}(\mu+L)}{\mu^{2}(1-\rho_{\mathcal{L}})^{4}}.$ (15)
The function $\bar{h}(T)$ grows linearly in $T$. Additionally, it comprises two terms: 1) the first term which captures the contribution from the centralized gradient descent optimization (see [58]) at each time instant without malicious agents, and 2) the following terms that include $\rho_{\mathcal{L}}$ which capture the contribution from distributing the optimization over a decentralized network (without malicious agents) that is characterized by the second largest eigenvalue modulus of $\overline{W}_{\mathcal{L}}$.
@@
$\lim_{t\rightarrow\infty}\|z_{i}(t)-x_{\mathcal{L}}^{\star}\|=0,\>\>\forall\>i\in\mathcal{L}.$
Moreover, if $\gamma(t)=\frac{2}{\mu(t+1)}$, then
$\frac{1}{|\mathcal{L}|}\sum_{i\in\mathcal{L}}\|z_{i}(T)-x_{\mathcal{L}}^{\star}\|^{2}\leq\min\left\{4\eta^{2},\frac{4\overline{h}(T)}{\mu T(T+1)}\right\},$ (16)
for any initial points $z_{i}(0)\in\mathcal{X}$, $i\in\mathcal{L}$, and any $T\geq 1$.

### v3

@@ Abstract
This work focuses on the problem of distributed optimization in multi-agent cyberphysical systems, where a legitimate agent’s iterates are influenced both by the values it receives from potentially malicious neighboring agents, and by its own self-serving target function. We develop a new algorithmic and analytical framework to achieve resilience for the class of problems where stochastic values of trust between agents exist and can be exploited. In this case, we show that convergence to the true global optimal point can be recovered, both in mean and almost surely, even in the presence of malicious agents. Furthermore, we provide expected convergence rate guarantees in the form of upper bounds on the expected squared distance to the optimal value. Finally, numerical results are presented that validate our analytical convergence guarantees even when the malicious agents compose the majority of agents in the network and where existing methods fail to converge to the optimal nominal points.
@@ 1 Introduction
We are interested in investigating the class of problems where the physicality of the system plays an important role in achieving new possibilities of resilience for these systems. Indeed, the physicality of cyberphysical systems has been shown to provide many new channels of verification and establishing inter-agent trust through watermarking [21], wireless signal characteristics [20, 22], side information [23], and camera or LiDAR data cross-validation [24]. By exploiting these physicality-based measurements, agents can extract additional information about the trustworthiness of their neighbors.
@@ 1.1 Related work
To harm the system, a malicious agent can send falsified data to their legitimate neighbors. If the legitimate agents are unaware that this data comes from malicious neighbors, then the malicious agents will succeed in controlling the system [34, 35, 36, 10, 11, 37, 38, 39]. To combat the harmful effect of an attack, the approach taken in [34, 35, 36] requires the pre-existence of a set of trusted agents such that all other agents (legitimate or malicious) are connected to at least one trusted agent. Nonetheless, this approach is unrealistic when communication is sporadic such as in robotic and ad-hoc networks. The approaches in [10, 11, 40, 37, 38, 39] rely on the agent data values to detect and discard malicious inputs. In general, data-based approaches have an upper bound on the number of tolerable malicious agents which cannot exceed half of the network connectivity; in some cases, this condition can be relaxed to half of the number of agents in the system - see, for example, [39]. Thus, they are not robust when malicious agents form a majority [37]. When the number of malicious agents exceeds the tolerable number, the attack succeeds and malicious agents evade detection. In contrast with the existing works, our proposed method provides a significantly stronger resilience to malicious activity by exploiting the physical aspect of the problem, i.e., the wireless medium. Thus, each legitimate agent can learn trustworthy neighbors while optimizing the system objective. Our prior work [41] studies the implications of the agents’ learning ability, with regards to the trustworthiness of their neighbors, on distributed consensus systems. This work considers the more general case of distributed optimization systems where the agent’s goal is to minimize the sum of their local objective function under limited information exchange.
@@ 1.2 Paper contribution
This work extends our earlier paper [1], which formalized the problem of resilient distributed optimization with trust observations and presented an upper bound on the convergence rate in the mean squared sense. This work excluded the proofs for these theoretical results due to space limitations. The key contributions of this work beyond those of [1] are as follows: 1) We present the analysis of all the proofs that were omitted from [1]. 2) We provide additional insights and discussions for the theoretical results that are included in [1]. Specifically, this work includes extensive discussions regarding the assumptions that are used in this paper regarding the statistics of the trust observations and their relaxations. 3) We extend the convergence in the second moment that follows from [1] and establish the almost sure convergence to the true optimal point. Additionally, we derive convergence in the $r$th mean for any $r>0$, both as a result of the almost sure convergence and by definition. These convergence results lead to an additional method to derive upper bounds on the convergence rate of the second moment which we present in this work. 4) We present additional numerical simulations that significantly extend the one-dimensional consensus with constraint setup to higher dimensional constrained optimization setups. Furthermore, we present in the supplementary material a comparison between the numerical simulations and the upper bounds we establish in this work.
@@ 2 Problem Formulation
We study the case where an unknown subset of the agents is malicious and the trustworthy agents are learning which neighbors they can trust. Thus, $\mathbb{V}=\mathcal{L}\cup\mathcal{M}$ where $\mathcal{L}$ is the nonempty and finite set of legitimate agents that execute computational tasks and share their data truthfully, while $\mathcal{M}$ denotes a finite set of agents that are not truthful. The sets $\mathcal{L}$ and $\mathcal{M}$ are defined solely for the sake of analysis, and none of the legitimate agents knows if it has malicious neighbors, or not, at any time. Throughout the paper, we will use the subscripts $\mathcal{L}$ and $\mathcal{M}$ to denote the various quantities related to legitimate and malicious agents, respectively.
@@ 2.1 Trust values
This model of inter-agent trust observations has been used in prior works [22, 41]. In general, the trust observations can be extracted from the physicality of the systems in various ways as elaborated in the recent survey [50]. For example, the works [51, 52] extract trust observation based on the characteristics of the received wireless signals that are used to transfer information between agents. The trust values then capture the likelihood of discrepancies between the agent’s reported location and its true location that is manifested in the signal characteristics, such as power and direction of arrival.
@@
(ii) [Independence of trust observations] The observations $\alpha_{ij}(t)$ are independent for all $t$ with $i\in\mathcal{L}$, $j\in\mathcal{N}_{i}$.
@@
Discussion: For the sake of simplicity of presentation, we assume that the expectation terms $E_{\mathcal{L}}$ and $E_{\mathcal{M}}$ are homogeneous (see Assumption 1(i)). We note that the closed-form analytical results derived in this work can be readily extended to the case where $E_{\mathcal{L}}\triangleq\min_{i\in\mathcal{L},j\in\mathcal{N}_{i}\cap\mathcal{L}}\liminf_{t\rightarrow\infty}\{\mathbf{E}[\alpha_{ij}(t)]-0.5\}$ and $E_{\mathcal{M}}\triangleq\max_{i\in\mathcal{L},j\in\mathcal{N}_{i}\cap\mathcal{M}}\mathop{\rm limsup}_{t\rightarrow\infty}\{\mathbf{E}[\alpha_{ij}(t)]-0.5\}$ are such that $E_{\mathcal{L}}>0$ and $E_{\mathcal{M}}<0$. Additionally, our analysis can be readily adapted to incorporate heterogeneity across clients and replace the uniform $E_{\mathcal{L}}$ and $E_{\mathcal{M}}$ terms with per link terms, i.e., $E_{ij}=\mathbf{E}[\alpha_{ij}(t)]-0.5$. Additionally, while in this work we derive a first-moment analysis by utilizing Hoeffding’s inequality, one can derive an analogous second-moment analysis by utilizing Bernstein’s inequality. Finally, we include in Section 4 a discussion regarding the relaxation of Assumption 1(ii) to include random processes with memory such as Markov chains and martingales.
@@ 2.2 The update rule of agents
The update rule for the legitimate agents
@@
To increase the convergence rate, we allow $\gamma(t)$ and $w_{ij}(t)$ to depend on an arbitrary parameter $T_{0}\geq 0$ which dictates how many trust observations a legitimate agent collects before starting to execute the distributed optimization protocol. We specify $\gamma(t)$ and $w_{ij}(t),\>i\in\mathcal{L}$, $j\in\mathcal{N}_{i}\cup\{i\}$, precisely later on in Section 3.
@@
Malicious agents $i\in\mathcal{M}$ choose values in the set $\mathcal{X}$ and they may or may not collaborate in choosing their values. We assume that their actions are not known, and thus we do not model them, but assume the worst-case scenario instead in our analysis. For simplicity of exposition, the dynamic (2.2) captures malicious inputs, where an adversarial agent $i\in\mathcal{M}$ sends all its legitimate neighbors identical copies of its chosen input $x_{i}(t)$ at time $t$. Let us denote by $x_{ij}(t)$ the input of a malicious agent $i$ to a legitimate agent $j$ at time $t$, then $x_{ij_{1}}(t)=x_{ij_{2}}(t)$ for every $j_{1},j_{2}\in\mathcal{N}_{i}\cap\mathcal{L}$. Nonetheless, our analytical results also hold for byzantine inputs where an adversarial agent $i\in\mathcal{M}$ can send its legitimate neighbors different inputs at time $t$. In this case $x_{ij_{1}}(t)$ need not be equal to $x_{ij_{2}}(t)$ for every $j_{1},j_{2}\in\mathcal{N}_{i}\cap\mathcal{L}$.
@@ 2.3 Assumptions on connectivity and objective functions
The set of legitimate agents is not empty, i.e., $\mathcal{L}\neq\emptyset$. Furthermore, the subgraph $\mathbb{G}_{\mathcal{L}}$ induced by the legitimate agents is connected.
@@ 3 Algorithm
Next, we present an algorithm that incorporates the legitimate agents’ learning of inter-agent trust values into the dynamic (2.2) through the choice of the time-dependent weights $w_{ij}(t)$. We recall that these weights depend on the parameter $T_{0}$. We utilize the parameter $T_{0}$ to enable faster convergence rates of the algorithm, nonetheless, $T_{0}$ is mostly used for analysis and it is not strictly necessary. In other words, as we show in Section 5, the algorithm converges to the optimal point $x_{\mathcal{L}}^{\star}$ for any choice of nonnegative integer $T_{0}$, including the special case where $T_{0}=0$. In this case, legitimate agents have no prior trust observations to rely on when they first decide whether to trust their neighbors.
@@ 3.1 The weight matrix sequence
$\beta_{ij}(t)=\sum_{k=0}^{t-1}\left(\alpha_{ij}(k)-0.5\right)\hbox{ \ for $t\geq 1,\>i\in\mathcal{L},\ j\in\mathcal{N}_{i}$},$ (4)
@@
We define the weight matrix $W(t)$ by choosing its entries $w_{ij}(t)$ as follows: for every $i\in\mathcal{L}$, $j\in\mathcal{L}\cup\mathcal{M}$
$w_{ij}(t)=\begin{cases}\frac{\mathds{1}_{\{t\geq T_{0}\}}}{2\cdot\max\{d_{i}(t),d_{j}(t)\}}&\text{if }j\in\mathcal{N}_{i}(t),\\ 0&\text{if }j\notin\mathcal{N}_{i}(t)\cup\{i\},\\ 1-\displaystyle{\sum_{m\in\mathcal{N}_{i}(t)}w_{im}(t)}&\text{if }j=i.\end{cases}$ (6)
@@
The dependence of the weights $w_{ij}(t)$ on the trust observation history $\beta_{ij}(t)$ comes in through the choice of time-dependent and random trusted neighborhood $\mathcal{N}_{i}(t)$ (see (5)). Consequently, some entries of the matrix $W(t)$ are also random, as seen from (6). The gradients $\nabla f_{i}(c_{i}(t))$ are stochastic due to the randomness of $c_{i}(t)$, however, they are not unbiased as typically assumed in stochastic approximation methods, including [53]. Thus, we cannot readily rely on prior analysis for stochastic approximation methods. However, as we show in our subsequent analysis, the variance of $\|\nabla f_{i}(c_{i}(t))\|$ decays sufficiently fast and allows convergence to the optimal point even in the presence of malicious agents.
@@ 4 Learning the sets of trusted neighbors
We can now conclude that there is a random but finite time $T_{f}$ such that there exists a legitimate agent $i$ which misclassifies the trustworthiness of at least one of its neighbors at time $T_{f}-1$ whenever $T_{f}\geq 1$, and all the legitimate agents classify the trustworthiness of their neighbors correctly at each time $t$ such that $t\geq T_{f}\geq 0$. We refer to the time $T_{f}$ as the “correct classification time”.
@@
$\beta_{ij}(t)$ $<0\text{ for all $t\geq T_{f}$ and all }i\in\mathcal{L},j\in\mathcal{N}_{i}\cap\mathcal{M}.$ (8)
Furthermore, if $T_{f}\geq 1$ there exists $i\in\mathcal{L}$ such that
@@
$\beta_{ij}(T_{f}-1)$ $\geq 0\text{ for some }j\in\mathcal{N}_{i}\cap\mathcal{M}.$ (9)
@@
We define $p_{c}(-1)$ in the above equation for the sake of notational completeness of Lemma 2. Furthermore, we define the following upper bound on the probability that a legitimate agent misclassifies one of its legitimate or malicious neighbors, in one of the times after observing $k$ trust values for each of its neighbors:
@@ 5 Asymptotic Convergence to Optimal Point
$\leq\gamma(t-T_{0})\|\nabla f_{i}(c_{i}(t))\|\leq\gamma(t-T_{0})G.\hskip 42.67912pt\qed$
Let us denote $d_{i,\mathcal{L}}\triangleq|\mathcal{N}_{i}\cap\mathcal{L}|+1$. Next, we define the doubly stochastic matrix $\overline{W}_{\mathcal{L}}\in[0,1]^{|\mathcal{L}|\times|\mathcal{L}|}$ with the entries $[\overline{W}_{\mathcal{L}}]_{i,j}$, for every $i,j\in\mathcal{L}$,
@@
Note that $\overline{W}_{\mathcal{L}}$ is the nominal weight matrix, i.e., the weight matrix that would be used in the absence of malicious agents. Let $\rho_{\mathcal{L}}=\sigma_{2}(\overline{W}_{\mathcal{L}})$ be the second largest singular value of $\overline{W}_{\mathcal{L}}$. Since $\mathbb{G}_{\mathcal{L}}$ is connected and $\overline{W}_{\mathcal{L}}$ is symmetric and doubly stochastic, the value $\rho_{\mathcal{L}}<1$ is equal to the second largest eigenvalue modulus of $\overline{W}_{\mathcal{L}}$. Additionally, our analysis holds for the lazy Metropolis weight as well, for which by [56, Lemma 2.2], the value $\rho_{\mathcal{L}}$ can be upper bounded by $(1-1/(71|\mathcal{L}|^{2}))$, while [57] improves the constant $1/71$ to $1/8$.
Next, define for $t\geq 0$ the following deterministic dynamic that excludes malicious agents:
@@
Let $\epsilon\in(0,1)$, $\mu_{\epsilon}\triangleq\mu(1-\epsilon)$, and denote for $T\geq 1$, the auxiliary error function which we utilize to characterize the convergence rate:
$\overline{h}_{\epsilon}(T)\triangleq\frac{G^{2}T}{\mu_{\epsilon}}+\frac{4G^{2}T}{\mu_{\epsilon}(1-\rho_{\mathcal{L}})}+\frac{2\eta G}{(1-\rho_{\mathcal{L}})^{2}}+\frac{4G^{2}}{\mu_{\epsilon}(1-\rho_{\mathcal{L}})^{3}}$
$+2(\mu_{\epsilon}/\epsilon+L)\bigg{[}\frac{\eta(\eta\mu_{\epsilon}+4G)}{\mu_{\epsilon}(1-\rho_{\mathcal{L}})^{2}}+\frac{4G^{2}\ln\left(\frac{T+2}{2}\right)}{\mu_{\epsilon}^{2}(1-\rho_{\mathcal{L}})^{2}}$
$\hskip 28.45274pt+\frac{4G(\eta\mu_{\epsilon}+G)}{\mu_{\epsilon}^{2}(1-\rho_{\mathcal{L}})^{3}}+\frac{G^{2}}{\mu_{\epsilon}^{2}(1-\rho_{\mathcal{L}})^{4}}\bigg{]}.$ (15)
The function $\bar{h}(T)$ grows linearly in $T$. Additionally, it comprises two terms: 1) the first term which captures the contribution from the centralized gradient descent optimization (see [58]) at each time instant without malicious agents (up to a scaling of $1/(1-\epsilon)$), and 2) the following terms that include $\rho_{\mathcal{L}}$ which capture the contribution from distributing the optimization over a decentralized network (without malicious agents) that is characterized by the second largest eigenvalue modulus of $\overline{W}_{\mathcal{L}}$.
@@
$\lim_{t\rightarrow\infty}\|z_{i}(t)-x_{\mathcal{L}}^{\star}\|=0\quad\hbox{for all }i\in\mathcal{L}.$
Moreover, if $\gamma(t)=\frac{2}{\mu_{\epsilon}(t+2)}$ where $\mu_{\epsilon}\triangleq\mu(1-\epsilon)$, then
$\frac{1}{|\mathcal{L}|}\sum_{i\in\mathcal{L}}\|z_{i}(T)-x_{\mathcal{L}}^{\star}\|^{2}\leq\min\left\{4\eta^{2},\frac{4\overline{h}_{\epsilon}(T)}{\mu_{\epsilon}T(T+1)}\right\},$ (16)
for any initial points $z_{i}(0)\in\mathcal{X}$, $i\in\mathcal{L}$, any $T\geq 1$, and any choice of $\epsilon\in(0,1)$.